OpenStudy (anonymous):

Help please. Fan and medal. Given the line 2x – 3y = 9 and the point (4, –1), find lines through the point that are (a) parallel to the given line and (b) perpendicular to it.

OpenStudy (anonymous):

Have to tried plotting the function?

OpenStudy (anonymous):

@Tommynaiter I still don't understand how to get the equations for the parallel and perpendicular lines.

OpenStudy (mrnood):

Do oyu know th estraight line equation y = mx+b ? rewrite your equation so it looks like that y = ??

OpenStudy (anonymous):

its called slope intercept form btw

OpenStudy (anonymous):

not estraight line equation

OpenStudy (anonymous):

I know the form the equation is supposed to be in. I don't know how to convert it

OpenStudy (anonymous):

all you need to do is isolate y

OpenStudy (anonymous):

so, subtract x from both sides, and divide both sides by y's coefficient

OpenStudy (mrnood):

@TheSavior you need to subtract 2x from each side then divide by 3

OpenStudy (mrnood):

correction divide by -3

OpenStudy (anonymous):

-3 IS y's coeffeicient lol

OpenStudy (anonymous):

Then y still wouldn't be isolated right?

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

oh never mind I see it now

OpenStudy (anonymous):

so now I have y=-3-2x right?

OpenStudy (anonymous):

No, move the x so that it is y=mx+b

OpenStudy (anonymous):

ok so y=2x-(-3) but what now

OpenStudy (anonymous):

and you didn't divide the -2 (x's coefficient also known as the slope, or m in this case) by -3

OpenStudy (anonymous):

that's still not right.

OpenStudy (anonymous):

well then idk

OpenStudy (anonymous):

when you move a number, you do not change whether it is positive or negative. The 2x should still be negative

OpenStudy (anonymous):

very close, it was y=-2x-3

OpenStudy (anonymous):

Now to answer A, you have a point

OpenStudy (anonymous):

in addition, you have the slope.

OpenStudy (anonymous):

Did you learn point-slope form?

OpenStudy (anonymous):

oh right, I forgot!

OpenStudy (mrnood):

y= 2/3 x -3

OpenStudy (anonymous):

I think she left

OpenStudy (anonymous):

now I'm just lost

OpenStudy (anonymous):

okay, it is still the same line, we just turned it into slope intercept form

OpenStudy (anonymous):

do you know point slope form?

OpenStudy (anonymous):

Kind of

OpenStudy (anonymous):

All you need to know is the formula

OpenStudy (anonymous):

since y-y1=m(x-x1), you only need a slope and a random point

OpenStudy (anonymous):

that is on the line

OpenStudy (anonymous):

you have a slope and a point, so you can plug in the y coordinate for y1, the x coordinate for x1.

OpenStudy (anonymous):

then plug in -2/3 for m

OpenStudy (anonymous):

that is parallel, because the slope is the same

OpenStudy (anonymous):

y-(-1)=-2/3(x-4)

OpenStudy (anonymous):

right or no?

OpenStudy (anonymous):

and what about the perpendicular line

OpenStudy (mrnood):

what do you know about slope of perpendicular lines?

OpenStudy (anonymous):

They have opposite slopes

OpenStudy (mrnood):

not opposite if one line has slope m then its perpendiclar has lope -1/m so what is the slope of the perp to your line?

OpenStudy (anonymous):

not a clue

OpenStudy (mrnood):

you need to go and review your text about 'point intercept' formula you need to know this before you can answer the question

OpenStudy (mrnood):

parallel lines have SAME slope perpendicular lines have slopes so that if one line has slope m then its perpendiclar has lope -1/m

OpenStudy (anonymous):

but how do you do that when the slope is -2/3

OpenStudy (mrnood):

your equatioon is y = 2/3 x -3 so the slope is 2/3 (NOT -2/3) you can see that from your plot - the slope is positive so what is -1/m if m=2/3

OpenStudy (anonymous):

@MrNood I don't know how to do that without getting some great big decimal